### v1

# Step-Aware Residual-Guided Diffusion for EEG Spatial Super-Resolution

Abstract
For real-world BCI applications, lightweight Electroencephalography (EEG) systems offer the best cost–deployment balance. However, such spatial sparsity of EEG limits spatial fidelity, hurting learning and introducing bias. EEG spatial super-resolution methods aim to recover high-density EEG signals from sparse measurements, yet is often hindered by distribution shift and signal distortion and thus reducing fidelity and usability for EEG analysis and visualization. To overcome these challenges, we introduce SRGDiff, a step-aware residual-guided diffusion model that formulates EEG spatial super-resolution as dynamic conditional generation. Our key idea is to learn a dynamic residual condition from the low-density input that predicts the step-wise temporal and spatial details to add and uses the evolving cue to steer the denoising process toward high density reconstructions. At each denoising step, the proposed residual condition is additively fused with the previous denoiser feature maps, then a step-dependent affine modulation scales and shifts the activation to produce the current features. This iterative procedure dynamically extracts step-wise temporal rhythms and spatial-topographic cues to steer high-density recovery and maintain a fidelity–consistency balance. We adopt a comprehensive evaluation protocol spanning signal-, feature-, and downstream-level metrics across SEED, SEED-IV, and Localize-MI and multiple upsampling scales. SRGDiff achieves consistent gains of up to 40% over strong baselines, proving its superiority in the task of EEG spatial super-resolution. Moreover, topographic visualizations comparison and substantial EEG-FID gains jointly indicate that our SR EEG mitigates the spatial–spectral shift between low- and high-density recordings. Our code is available at https://anonymous.4open.science/r/DDPM-VAE-6F6B.
1 Introduction
Electroencephalography (EEG) is a noninvasive technique for monitoring the brain’s electrical activity, with widespread applications in neuroscience and clinical practice—ranging from brain–computer interfaces and epilepsy diagnosis to emotion recognition Jiang et al. (2025). However, EEG’s spatial resolution is inherently constrained by the number of scalp electrodes and the volume‐conduction effect Li et al. (2025a). High‐density (HD) systems with hundreds of channels can mitigate these issues but are costly, cumbersome to deploy, and uncomfortable for extended wear, whereas low‐density (LD) setups (e.g., 8 or 16 electrodes) are far more practical yet suffer from severe under‐sampling bias Wang et al. (2025). Indeed, as illustrated in Figure 1(c), the inter‐channel activation patterns of 256‐channel HD EEG diverge dramatically from those of 16‐channel LD EEG, highlighting the strong bias in sparse recordings. EEG spatial super-resolution (SR) has therefore garnered growing attention, with methods that reconstruct high-density EEG from sparse recordings increasingly explored and applied.
Traditionally, EEG spatial super‐resolution has relied on direct feature‐mapping techniques that learn an end-to-end mapping from low-density to high-density representations. These methods fall into two main categories: one employs convolutional neural networks or Transformers to upsample LD feature maps into HD ones Tang et al. (2022), and the other leverages generative adversarial networks-based architectures that synthesize SR EEG signals conditioned on LD inputs Wang et al. (2024). However, by treating the mapping as a static projection, these approaches often oversimplify the complex, nonlinear inter-electrode dependencies or demand vast amounts of training data and compute, resulting in overly smooth, detail-poor reconstructions that fail to capture true spatial consistency. Figure 1(c) indicates that such feature-mapping methods merely extend LD information, rather than recovering authentic HD channel relationships.
Figure 1: (a) Existing static guidance strategy vs. (b) our residual guidance strategy for EEG super-resolution, and (c) corresponding topographical maps of LD input, ESTformer output, GT HD EEG, and SRGDiff reconstruction.
Recently, diffusion models have been widely applied to time-series generation and missing-data imputation Huang et al. (2025); Yuan & Qiao (2024); Li et al. (2025b). In this context, EEG spatial super-resolution can be cast as conditional generation, where LD observations guide the recovery of HD signals. Within this line of work, researchers mainly focused on conditioning strategies through concatenating low-density features with the noise input Vetter et al. (2024) or using cross-attention between modalities Wang et al. (2025) as shown in Figure 1(a). While effective in practice, these approaches remain susceptible to a consistency–fidelity trade-off. Interpolation-oriented SR tends to cause distribution shift, making reconstructions adhere too closely to the LD observation and deviate from the HD ground truth. Conversely, generation-oriented SR often introduces distortion, producing HD-like content that fails to remain consistent with the LD input.
To tackle these challenges, we introduce Step-aware Residual-Guided Diffusion (SRGDiff) for EEG Spatial Super-Resolution, which reframes super-resolution as a dynamic conditional generation task. The core idea is to estimate the forward-noising residual from low-density channels, and use it as a per-step corrective direction in the reverse process. Technically, SRGDiff first encodes the low-density EEG with a pre-trained VAE encoder to obtain a compact latent and multi-scale features, and applies forward diffusion to the high-density latent. At each reverse step, a lightweight residual head predicts a path residual from the low-density features and uses it as a directional correction that is additively fused with the previous denoising features to form an incremental feature. The feature is then weighted with a step-dependent affine modulation estimated from the low-density features and the timestep embedding, yielding the current denoised features. This loop repeats over timesteps, coupling the low-density forward-noising and high-density reverse-denoising trajectories and progressively steering samples toward the high-density manifold. Our main contributions in this work can be summarized as follows:
  • We recast EEG spatial super-resolution as dynamic conditional generation, coupling the LD forward–noising trajectory with the HD reverse–denoising trajectory to balance consistency with the LD observation and fidelity to the HD target.
  • We propose a dynamic residual guidance paradigm: the path residual estimated from LD inputs serves as a per-step directional correction and is fused additively for incremental updates, yielding a stable, step-aware sampling scheme transferable to general SR/MDI settings.
  • We establish a three-level evaluation protocol across three datasets, covering signal-level (temporal consistency, spectral fidelity, spatial topology), feature-level (representation quality), and downstream-level (classification accuracy), which provides a comprehensive assessment beyond pointwise error.
2 Related Work
Diffusion Models for Missing Data Imputation.
Diffusion‐based models have emerged as a powerful framework for time‐series imputation, leveraging denoising diffusion processes to reconstruct missing values. Diffusion-TS Yuan & Qiao (2024) demonstrates interpretable conditional generation across diverse time‐series without domain‐specific priors. More recently, RDPI Liu et al. (2025) further enhances precision and efficiency by first generating coarse estimates of missing values through deterministic interpolation, then conditioning a diffusion model on both observed data and these estimates to iteratively refine residual errors. SaSDim Zhang et al. (2024) introduces self‑adaptive noise scaling to preserve spatial dependencies within sensor networks, while SADI Islam et al. (2025) integrate self‑attention mechanisms to handle partial data missing.
EEG Spatial Super-Resolution.
Early attempts at EEG spatial super‐resolution adapted image‐based frameworks to reconstruct dense electrode maps from sparse recordings. EEGSR‐GAN Corley & Huang (2018) first applied adversarial training to hallucinate missing channels. ESTformer Li et al. (2025a) then introduced spatiotemporal transformers to model long‐range dependencies across electrodes and successfully capture global patterns. More recent diffusion‐based and attention‐driven approaches have sought to address these limitations. DDPM‐EEGVetter et al. (2024) leverages denoising diffusion probabilistic models to iteratively refine spatial patterns. STAD Wang et al. (2025) tackles this by decomposing spatial–temporal interactions into spatial-temporal attention streams. This diffusion-based generative paradigm improves diversity and spectral fidelity compared to GANs, yet their static condition can still lead to distribution drifts and distortion.
Residual diffusion in related domains.
Several recent works incorporate residual signals into diffusion models. Ou et al. (2024) synthesize PET from MRI by learning a modality residual under prior information, i.e., a static cross-modal gap that conditions generation. Zhu et al. (2024) reconstruct event-driven video by predicting temporal residuals with inter-frame differences as the generation target to recover dynamics. Mao et al. (2025) address medical segmentation by learning a residual-to-prior that corrects a coarse segmentation, improving calibration and efficiency. These designs either treat the residual as a fixed target/offset or inject it once as a global prior, with weak coupling to the step-by-step reverse dynamics. In contrast, our method targets EEG spatial super-resolution and introduces a dynamic, step-aware residual direction that is re-estimated at every reverse step from the LD observation, and the timestep embedding.
Figure 2: SRGDiff overview. (a) Overall architecture: Low-density EEG $X^{L}$ conditions the latent reverse process. RDM predicts a residual direction from $X^{L}$ and current decoder features. SMM provides step-aware affine parameters to fuse the residual and modulate activations. (b) Residual learning: At each step, the predicted residual guides denoising, and the residual derived from the forward noising process provides supervision via a residual loss.
3 Preliminaries and Dynamic Conditional Formulation
Data and Latent Space. Let $X^{L}\!\in\!\mathbb{R}^{C_{L}\times Length}$ and $X^{H}\!\in\!\mathbb{R}^{C_{H}\times Length}$ denote low- and high-density EEG with $C_{H}\!>\!C_{L}$. A pre-trained VAE encoder $E$ maps signals to a latent space $z=E(X^{H})$ and a feature extractor $F$ provides LD features as condition $c=F(X^{L})$ to condition generation. In practice, we reuse the VAE encoder as the feature extractor to obtain condition $c$.
Forward Diffusion on the HD Latent. We corrupt the HD latent with a standard Gaussian forward process
$q(z_{t}\mid z_{z-1})=\mathcal{N}\!\Big(\sqrt{\alpha_{t}}\,z_{t-1},\,(1-\alpha_{t})I\Big).\qquad t=1,\dots,T$ (1)
Dynamic Conditional Reverse Process. In the reverse generation stage, sampling begins from an isotropic Gaussian noise initialization $\hat{z}_{T}\sim\mathcal{N}(0,I)$, and the model iteratively predicts $\hat{z}_{t-1}$ from $\hat{z}_{t}$ until recovering the final latent representation $\hat{z}_{0}$. Unlike conventional diffusion models that rely solely on a base denoiser, SRGDiff explicitly conditions each reverse step on low-density EEG observations, thereby coupling the LD forward-noising trajectory with the HD reverse-denoising trajectory. The reverse denoiser is defined as:
$p_{\theta}(\hat{z}_{t-1}\mid\hat{z}_{t},c)=\mathcal{N}\!\Big(\underbrace{\mu_{\theta}(\hat{z}_{t},c)}_{\text{base denoiser}}+\underbrace{(\gamma_{t},\beta_{t},r_{\phi}(c,t))}_{\text{dynamic conditional update}},\,\beta_{t}I\Big).$ (2)
Here, the base denoiser $\mu_{\theta}(\hat{z}_{t},c)$ is implemented as a U-Net that estimates the noise component. To further enhance temporal fidelity and spatial coherence, we augment the base denoiser with two lightweight modules that inject step-wise conditional guidance from LD features:
  • Residual Direction Module (RDM). At each timestep, RDM predicts a path residual $r_{\phi}(c,t)$ from the LD features and applies it as a directional correction:
    $\hat{z}^{RDM}_{t-1}=\hat{z}_{t}+r_{\phi}(c,t).$ (3)
  • Step-Aware Modulation Module (SMM). SMM calibrates the residual update with timestep-aware affine modulation. Specifically, it predicts a scale $\gamma_{t}$ and bias $\beta_{t}$ from LD features and the timestep embedding, and applies them to the residual-corrected state:
    $\hat{z}^{SMM}_{t-1}=\gamma_{t}\odot\hat{z}^{RDM}_{t-1}+\beta_{t}.$ (4)
Together, the pair $(r_{\phi},\gamma_{t},\beta_{t})$ realizes dynamic conditional generation: at every denoising step, LD features provide both a directional residual and a step-dependent modulation strength, yielding a stable and temporally consistent correction of the reverse diffusion process.
4 Proposed Method
This section presents the step-aware residual-guided diffusion framework, and outlines its architecture and core components. SRGDiff reframes SR as dynamic conditional generation that couples the LD forward-noising trajectory with the HD reverse-denoising trajectory, using an LD-estimated residual as a per-step corrective direction and a step-aware calibration to modulate its strength. The framework comprises four parts: the latent diffusion model backbone, Residual Direction Module (RDM) for additive directional updates, Step-Aware Modulation Module (SMM) for step-dependent modulation, and the overall training strategy. An overview is provided in Figure 2, and the following subsections describe each component in detail.
4.1 Latent Diffusion Model Backbone
Our backbone consists of a VAE that builds the latent space and a denoising U-Net that performs diffusion in that space. The VAE follows the EEG autoencoding setup of Aristimunha et al. (2023). We train an encoder–decoder $(E,D)$ on HD EEG $X^{H}$ to obtain $z=E(X^{H})$ $\widehat{X}^{H}=D(\widehat{z}), and$ optimize a reconstruction–regularization objective
$\mathcal{L}_{\text{VAE}}=\|\widehat{X}^{H}-X^{H}\|_{2}^{2}+\lambda_{\text{spec}}\big\|\mathrm{STFT}(\widehat{X}^{H})-\mathrm{STFT}(X^{H})\big\|_{1}+\lambda_{KL}\,\mathrm{KL}\!\left(q_{E}(z\!\mid\!X^{H})\,\|\,\mathcal{N}(0,I)\right).$ (5)
Empirically, we set the spectral weight to $0.1$ and the KL weight to $10^{-4}$. After convergence, $(E,D)$ are frozen. On top of this latent space, we adopt a latent-diffusion model in the style of Rombach et al. (2022).
4.2 Residual Direction Module
In the context of EEG spatial SR, most diffusion approaches condition the U-Net via feature concatenation or cross-attention. We instead turn EEG spatial SR into finer and step-aware conditioning by learning residual direction from low-density recordings. We use the VAE encoder to extract multi-scale condition $c$ and an RDM head $R_{\phi}$ takes $(c,\,\tau(t))$ to predict a residual $Res_{t}$ in the encoder feature space, which acts as a per-step directional correction to the reverse process. Concretely, we first sample a timestep $t$, encode the HD EEG to obtain the latent $z_{0}=E(X^{H})$, and draw
$z_{t}\;=\;\sqrt{\bar{\alpha}_{t}}\,z_{0}\;+\;\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\qquad\epsilon\sim\mathcal{N}(0,I).$ (6)
In the forward process, we obtain at each step the noise-corrupted latent of the HD EEG and use this sequence of step-dependent features as supervision targets for the residual. The residual labels span $t=0,\dots,T$ and are defined as $\delta z_{t}\;:=\;z_{0}\;-\;z_{t},$.
In the reverse process, we estimate $\delta z_{t}$ from the low-density EEG to supply the step-wise temporal and spatial details required for denoising. We introduce a lightweight convolutional predictor $R_{\phi}$ that takes the timestep embedding $\tau(t)$ and LD features $c=F(X^{L})$ as input and outputs the residual feature $Res_{t}$, and trained by
$Res_{t}\;=\;R_{\phi}\!\big(\tau(t),\,c\big),\quad\mathcal{L}_{\mathrm{res}}\;=\;\sum_{t=0}^{T}\big\|\,Res_{t}-\delta z_{t}\big\|_{2}^{2}.$ (7)
Finally, the predicted residual feature is then added to $\hat{z}_{t}$ as an incremental update:
$\hat{z}_{t}^{RDM}\;=\;LayerNorm(\hat{z}_{t})\;+\;Res_{t}.$ (8)
4.3 Step-Aware Modulation Module
After obtaining $\hat{z}_{t}^{\mathrm{RDM}}$, we further modulate the current step to control the extent to which the residual condition influences denoising. To enforce temporal fidelity, SMM explicitly weighted the current diffusion timestep with a step-dependent affine modulation estimated from the low-density features and the current timestep embedding.
Specifically, SMM first encodes the low‐density EEG through a lightweight 1D convolutional network $E_{SMM}$ to produce a feature map $h_{t}$. To enable the conditioning to recognize the current diffusion step, SMM maps each sampled timestep $t$ into a sinusoidal time embedding $e_{t}$. A learnable weight $\sigma_{t}$ that decays linearly with $t$ balances these two streams, yielding a fused feature:
$\widetilde{h}_{t}=\sigma_{t}h_{t}+(1-\sigma_{t})e_{t}=\gamma_{t}E_{SMM}(c)+(1-\gamma_{t})e_{t}.$ (9)
For spatial coherence, we adopt an affine calibration mechanism. The fused feature $\widetilde{h}_{t}$ is passed through two MLPs $MLP_{\gamma}$ and $MLP_{\beta}$ to predict channel‐wise scale $\gamma^{c}_{t}$ and bias $\beta^{c}_{t}$:
$\hat{z}_{t}^{SMM}=\gamma_{t}\,\odot\,\hat{z}_{t}^{RDM}\;+\;\beta^{c}_{t}=MLP_{\gamma}(\widetilde{h}_{t},t)\,\odot\,\hat{z}_{t}^{RDM}\;+\;MLP_{\beta}(\widetilde{h}_{t},t).$ (10)
Finally, SRGDiff feeds the updated latent $\hat{z}_{t}$ into the U-Net decoder to obtain the next denoised state $\hat{z}_{t-1}$.
4.4 Training Strategy
To stabilize optimization and decouple latent representation learning from conditional diffusion modeling, we adopt a two-stage training strategy.
Stage 1: VAE Pre-training. We first train the VAE encoder-decoder pair on high-density EEG data to obtain a stable and structured latent space. After convergence, the VAE parameters are frozen to provide fixed latent representations for subsequent diffusion modeling.
Stage 2: Residual-Guided Latent Diffusion. On the frozen latent space, the final training objective is a weighted combination of the three terms:
$\mathcal{L}_{\text{Stage 2}}=\mathbb{E}_{z_{0},\epsilon,t}\big[\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|_{2}^{2}\big]+\lambda_{res}\sum_{t=1}^{T}\|R_{\varphi}(c,t)-(z_{t}-z_{0})\|_{2}^{2}+\lambda_{SMM}(\|\gamma_{t}-1\|_{2}^{2}+\|\beta_{t}\|_{2}^{2}).$ (11)
Empirically, we set the residual weight to $1$ and the SMM weight to $10^{-2}$. The term $\lambda_{\text{SMM}}\big(\|\gamma_{t}-1\|_{2}^{2}+\|\beta_{t}\|_{2}^{2}\big)$ serves as a regularization component to prevent excessively large values of $\gamma_{t}$ and $\beta_{t}$, thereby stabilizing the training dynamics.
5 Experiments
5.1 Downstream Datasets
In this study, we employ three publicly available EEG datasets. The SEED dataset Zheng & Lu (2015) uses $15$ film clips of approximately four minutes each as emotional stimuli to induce stable and continuous emotional responses in three categories: positive, neutral, and negative. Data were acquired via $62$ channels at $1000$ Hz, downsampled to $200$ Hz, band-pass filtered (07575 Hz) and segments with faulty sensors removed. SEED-IV Zheng et al. (2018) extends SEED by using the same $15$ subjects and $62$-channel setup ($1000$ Hz) but adds music and image stimuli to evoke happiness, sadness, fear and neutral states; preprocessing mirrors that of SEED. Localize-MI Mikulan et al. (2020) contains 61 presurgical sessions from seven drug-resistant epilepsy patients, where $256$-channel scalp EEG was recorded at $8000$ Hz during 0.10.155 mA intracerebral single-pulse stimulation; preprocessing includes a $0.1$ Hz high-pass filter, notch filter, bad-channel/trial removal and alignment of trials to the -$300$ ms to +$50$ ms stimulus-artifact window.
5.2 Experiment Setup
Data Preprocessing.
The experimental setup for the EEG super-resolution task follows the ESTformer and STAD frameworks. The preprocessed EEG signals were segmented into fixed-length windows: continuous, non-overlapping $4$-second windows for SEED and SEED-IV datasets, and $260$ ms windows (from $250$ ms before stimulation to $10$ ms after) for Localize-MI. In SEED and SEED-IV, we designed different super-resolution scale factors ($2\times$, $4\times$ and $8\times$) to evaluate reconstruction performance. The selection of visible channels and the super-resolution scaling factors follow the configurations used in ESTformer. For Localize-MI, due to the high channel density, we applied more extensive scale factors ($2\times$, $4\times$, $8\times$, $16\times$).
Training & Environment Settings.
For each dataset, we split the data into train/test with an $80\%/20\%$ ratio and reserve $10\%$ of the training portion as a validation set. Stage I is trained only on the HD signals from the training split. Stage II is trained on paired LD/HD samples constructed from the same training split by masking HD channels according to the target SR scale; the validation set is used for early stopping and hyperparameter selection. The held-out test split is used once for final reporting, with no fine-tuning.
Baselines.
We compare SRGDiff with strong EEG SR and time-series imputation baselines: ESTformer Li et al. (2025a) and STAD Wang et al. (2025) (transformer-/diffusion-based EEG SR), DDPMEEG Vetter et al. (2024) (diffusion for ECoG SR), SaSDim Zhang et al. (2024) and SADI Islam et al. (2025) (advanced missing data imputation), and the two-stage residual method RDPI Liu et al. (2025). We use authors’ official implementations when available and otherwise provide carefully verified reimplementations, applying their recommended hyperparameters and unifying training epochs and sampling steps across methods.
Model Ref Metric 2 4 8 16
[…]
Table 1: Performance of all methods on Localize-MI across different channel settings.
Evaluation Protocol.
[…]
5.3 Main Results
We report signal-level reconstruction quality in Tables 1 and 2. For the most demanding $16\times$ upsampling on Localize-MI, SRGDiff attains an NMSE of $0.3457$, over $15\%$ lower than DDPMEEG’s $0.4076$, indicating that dynamic conditioning effectively guides the diffusion model to generate super-resolved signals that closely approximate the true high-density data. Its PCC improves from $0.6739$ to $0.7322$, and its SNR increases from $3.27$ dB to $4.02$ dB (over $22\%$), demonstrating that under challenging settings the dynamic conditioning still learns the HD trend while maintaining a favorable signal-to-noise ratio. On SEED with high temporal variability and frequent outliers, SRGDiff reduces the $2\times$ NMSE from ESTformer’s $0.3288$ to $0.1632$ (a reduction of more than $50\%$) and raises PCC from $0.8368$ to $0.9102$. A similar pattern is observed on SEED-IV, where NMSE drops to $0.1663$ versus $0.3448$ and PCC increases to $0.9113$ versus $0.8106$, indicating that despite lower SNR, the dynamic conditioning exhibits strong generalization.
Figure 3: EEG-FID evaluation across three datasets compared with ESTformer and STAD.
We also report EEG-FID results in Figure 3, and our method consistently achieves the lowest FID scores across SEED, SEED-IV, and Localize-MI datasets under different scale factors. These results indicate that our approach generates EEG signals that are statistically closer to the real distribution in the temporal domain.
Model Metric SEED (62) SEED-IV (62)
2 4 8 2 4 8
[…]
Table 2: Performance comparison of different models on SEED and SEED-IV datasets across different channel settings.
5.4 Topology Visualization Results
We further analyze the spectral fidelity of generated signals by visualizing EEG topographic maps under different scale factors. An EEG topographic map projects the power spectral density (PSD) of each channel onto the scalp surface, providing an intuitive representation of the spatial distribution of oscillatory energy. As shown in Figure 4, although our reconstructed signals still exhibit minor differences from the original data, they preserve a high degree of overlap in critical regions with strong PSD responses.
Figure 4: Visualization of EEG topographic maps between ground-truth and reconstructed EEG signals by ESTformer, STAD and SRGDiff.
5.5 Downstream tasks
Method SEED SEED-IV MI
2 4 8 2 4 8 2 4 8 16
GT 0.7152 0.7152 0.7152 0.7027 0.7027 0.7027 0.8368 0.8368 0.8368 0.8368
LR 0.4981 0.4702 0.4424 0.5685 0.5618 0.5011 0.7208 0.6534 0.5219 0.3862
SaSDim 0.5097 0.4793 0.4429 0.5794 0.5692 0.4912 0.7193 0.6519 0.5237 0.3845
SADI 0.5137 0.4834 0.4456 0.5718 0.5644 0.4987 0.7215 0.6634 0.5314 0.3957
RDPI 0.5044 0.4802 0.4531 0.5591 0.5741 0.5071 0.7230 0.6624 0.5210 0.3892
DDPMEEG 0.4738 0.4610 0.4238 0.5548 0.5487 0.4838 0.7015 0.6387 0.5187 0.3767
ESTformer 0.6887 0.6509 0.6057 0.6782 0.6500 0.5084 0.7445 0.6033 0.4739 0.4391
STAD 0.5437 0.5249 0.4610 0.6651 0.6410 0.4977 0.7589 0.6797 0.6344 0.5384
SRGDiff 0.7019 0.6812 0.6273 0.6821 0.6558 0.5127 0.7641 0.7163 0.6806 0.5887
Table 3: Classification accuracy comparison across different methods and datasets. GT represents the ground truth performance. Best results are shown in bold, second-best are underlined.
Table 3 reports results on the three datasets under various super-resolution scales. As the scale grows, accuracy for both raw and super-resolved inputs declines, yet SRGDiff’s reconstructions consistently maintain a clear advantage. In particular, SRGDiff greatly outperforms missing-value imputation methods and leads all other spatial imputation approaches. At $2\times$ scale factor, its classification accuracy approaches that obtained from the original full-channel recordings. We further compared the runtime efficiency of different methods as shown in Table 10. Although our proposed SRGDiff is slightly slower than the transformer-based ESTformer, it can still complete EEG super-resolution within 0.1s, which meets the real-time requirement in practical applications. The detailed runtime statistics of all models are provided in the Appendix.
5.6 Ablation Study
To evaluate the contribution of each module in SRGDiff to EEG super-resolution reconstruction, we conducted ablation studies comparing SRGDiff with three variant models. LDM retains only the pre-trained VAE and DDIM; LDM+SMM preserves the Step-aware modulation module; LDM+RDM retains the Residual Direction Module. All models were tested under the same experimental settings.
Figure 5: Ablation study performance comparison between SRGDiff and three variant models on the SEED dataset.
Figure 5 reports NMSE and PCC across $2\times$, $4\times$ and $8\times$ upsampling in SEED dataset. At $8\times$, adding SMM to the baseline LDM cuts NMSE from $0.86$ to $0.45$ with 4747% reduction and boosts PCC from $0.34$ to $0.69$, demonstrating its effectiveness in temporally aligning the denoising trajectory. Incorporating RDM yields a comparable NMSE reduction with 4444% and raises PCC to $0.67$, highlighting its role in injecting prior information for spatial consistency. When combined in SRGDiff, these modules further decrease NMSE to $0.34$ with 6060% overall reduction and elevate PCC to $0.81$. More ablation results in SEED-IV and Localize-MI datasets are shown in Figure 10 in the Appendix.
6 Conclusion
We introduced SRGDiff, a step-aware residual-guided diffusion model that reframes EEG spatial super-resolution as guided HD generation with a step-aware residual direction and adaptive modulation. Across SEED, SEED-IV, and Localize-MI, SRGDiff consistently improves signal-level metrics, achieves the best EEG-FID across scales, and better preserves spectral and scalp topographies. Downstream evaluations further show higher accuracy on emotion recognition and patient-wise classification, indicating that the reconstructed signals are not only visually and statistically closer to HD EEG but also more useful for analysis. These results validate that explicit, step-wise conditioning on sparse inputs is both necessary and effective for high-fidelity, topology-preserving EEG super-resolution.
7 Ethics Statement
This work adheres to the ICLR Code of Ethics. In this study, no human subjects or animal experimentation was involved. All datasets used, including SEED, SEED-IV and Localize-MI, were sourced in compliance with relevant usage guidelines, ensuring no violation of privacy. Details could be found in Appendix. We have taken care to avoid any biases or discriminatory outcomes in our research process. No personally identifiable information was used, and no experiments were conducted that could raise privacy or security concerns. We are committed to maintaining transparency and integrity throughout the research process.
8 Reproducibility Statement
We have made every effort to ensure that the results presented in this paper are reproducible. All code and datasets have been made publicly available in an anonymous repository to facilitate replication and verification. The experimental setup, including training steps, model configurations, and hardware details, is described in detail in the paper. We have also provided a full description of SRGDiff, to assist others in reproducing our experiments.
We believe these measures will enable other researchers to reproduce our work and further advance the field.
Appendix A LLM Usage
Large Language Models (LLMs) were used to aid in the writing and polishing of the manuscript. Specifically, we used an LLM to assist in refining the language, improving readability, and ensuring clarity in various sections of the paper. The model helped with tasks such as sentence rephrasing, grammar checking, and enhancing the overall flow of the text.
It is important to note that the LLM was not involved in the ideation, research methodology, or experimental design. All research concepts, ideas, and analyses were developed and conducted by the authors. The contributions of the LLM were solely focused on improving the linguistic quality of the paper, with no involvement in the scientific content or data analysis.
The authors take full responsibility for the content of the manuscript, including any text generated or polished by the LLM. We have ensured that the LLM-generated text adheres to ethical guidelines and does not contribute to plagiarism or scientific misconduct.
Appendix B Dataset Preprocess Details
B.1 Dataset Details
We use three publicly available EEG datasets: SEED, SEED-IV, and Localize-MI.
SEED (available at http://bcmi.sjtu.edu.cn/~seed/) consists of recordings from $15$ subjects watching emotion-eliciting film clips ($\approx$ 44 min each) designed to induce positive, neutral, and negative states. Data were acquired with a $62$-channel 102010–20 montage at $1000$ Hz, downsampled to $200$ Hz, and band-pass filtered to 0.5750.5–75 Hz. We manually inspected and removed sessions with sensor faults as depicted in Figure 6 and Figure 7.
Figure 6: Example of a raw SEED EEG segment with sensor faults highlighted.
Figure 7: Example of a raw SEED EEG segment with sensor faults highlighted.
SEED-IV (available at http://bcmi.sjtu.edu.cn/~seed/seed-iv.html) is a publicly available EEG dataset designed for emotion recognition research. It includes four emotional categories: happiness, sadness, neutrality, and fear. Emotional states are elicited using two types of stimuli: music and images. EEG recordings were collected from 15 subjects using a 62-channel system with a sampling rate of 1000 Hz.During preprocessing, the EEG signals were downsampled to 200 Hz and filtered with a bandpass filter ranging from 0.5 to 75 Hz. To ensure data quality, visually corrupted or invalid trials were manually excluded.
Localize-MI (available at https://doi.org/10.12751/g-node.1cc1ae) is a high-density intracranial EEG dataset from seven drug-resistant epilepsy patients during $61$ presurgical sessions. Stereo–EEG electrodes delivered single-pulse biphasic currents (0.150.1–5 mA), and $256$ channels were recorded at $8000$ Hz. Preprocessing included $0.1$ Hz high-pass filtering, notch filters at 5050/100100/150150/$200$ Hz, bad-channel/trial removal, and trial alignment using stimulation artifact peaks (300–300 to +50+50 ms window). In the Localize-MI dataset, we designed a binary classification task (epileptic vs. nonepileptic) to evaluate the effectiveness of synthetic super-resolution EEG (SR EEG) in detecting epileptic abnormalities. Specifically, EEG signals recorded before electrical stimulation are labeled as nonepileptic, while those recorded during stimulation are labeled as epileptic. The detailed experimental setup follows the description provided in the STAD Wang et al. (2025) model section.
B.2 More Experimental Details
We follow ESTformer and STAD slicing strategies. Preprocessed signals are windowed into fixed lengths: SEED and SEED-IV use non-overlapping $4$ s segments, while Localize-MI retains 250–250 ms to +10+10 ms around each stimulus ($260$ ms total). We randomly split 8080% for training and 2020% for testing, yielding $24265\times 62\times 800$ train / $6067\times 62\times 800$ test samples for SEED; $29199\times 62\times 800$ train / $7300\times 62\times 800$ test for SEED-IV; and $1914\times 256\times 2081$ train / $479\times 256\times 2081$ test for Localize-MI. For SEED and SEED-IV we evaluate $2\times$, $4\times$, and $8\times$ super-resolution; for Localize-MI we additionally include $16\times$. As shown in Figure 8, Localize-MI employs a 256-channel intracranial grid, while Figure 9 shows the 62-channel scalp montage used in SEED-IV and SEED.
Figure 8: Electrode topology of the Localize-MI dataset (256 intracranial channels).
Figure 9: Electrode topology of the SEED-IV and SEED datasets (62 scalp channels).
Appendix C More SRGDiff Model Details
C.1 Variational Autoencoder
In this paper, Variational Autoencoder (VAE) follows the AutoencoderKL design Aristimunha et al. (2023), comprising a convolutional encoder, a latent distribution (mean and variance) with KL regularization toward $\mathcal{N}(0,I)$, and a decoder with deconvolutions and upsampling. We augment both encoder and decoder with attention layers (multi-head and non-local attention), residual connections, and GroupNorm to capture global EEG features while ensuring stable training and efficient latent representations.
C.2 DDIMScheduler
The DDIMScheduler manages noise scheduling and sampling in the forward and reverse diffusion processes. It supports multiple noise prediction types and variance strategies. At each step, it computes the noise coefficient, predicts the denoised sample, clips values for numerical stability, and injects random perturbations to control output diversity.
Appendix D Parameter Study
D.1 VAE Latent Shape Selection
We found that the latent shape balances reconstruction precision and generalization. Higher dimensions capture more detail but risk overfitting, while lower dimensions blur outputs. We experimented across the three datasets and selected a latent of $32\times 400$ for SEED/SEED-IV and $64\times 500$ for Localize-MI, which yielded optimal NMSE, PCC, and SNR. Table 4, Table 5 and Table 6 shows the performance of SRGDiff in different latent shapes.
Shape NMSE PCC SNR (dB)
$64\times 400$ 0.15 0.93 7.75
[…]
$16\times 200$ 0.20 0.89 6.81
$8\times 400$ 0.16 0.92 7.26
Table 4: SEED: VAE latent shape selection results
Shape NMSE PCC SNR (dB)
$64\times 400$ 0.16 0.92 7.58
[…]
$16\times 200$ 0.21 0.87 6.86
$8\times 400$ 0.19 0.91 7.23
Table 5: SEED-IV: VAE latent shape selection results
Shape NMSE PCC SNR (dB)
$128\times 1000$ 0.13 0.94 8.60
[…]
$32\times 500$ 0.15 0.92 7.61
$32\times 1000$ 0.13 0.94 8.62
Table 6: Localize-MI: VAE latent shape selection results
D.2 Diffusion Hyperparameters
D.2.1 Diffusion Schedules
We compare linear and cosine noise schedules. Linear adds noise at a constant rate but may cause instability at endpoints; cosine offers smoother transitions and better performance for long diffusion chains. We fixed $1000$ timesteps with cosine scheduling and evaluated NMSE, PCC, and SNR on the latent reconstructions to choose this setting as shown in Table 7.
Dataset Schedule NMSE PCC SNR (dB)
SEED Linear 0.42 0.71 4.18
Cosine 0.20 0.86 7.15
SEED-IV Linear 0.51 0.66 4.02
Cosine 0.19 0.88 7.24
Localize-MI Linear 0.14 0.93 8.39
Cosine 0.11 0.95 8.88
Table 7: Comparison of noise schedules on three datasets (NMSE, PCC, SNR).
D.2.2 Training Timestep Lengths
We also tested different training timestep lengths ($200$, $1000$, $2000$). Larger values introduce stronger noise but make denoising harder; smaller values lack coverage of high-noise regimes. Using cosine scheduling, the results in Table 8 exhibit that $1000$ timesteps to be optimal across datasets.
Dataset Steps NMSE PCC SNR (dB)
SEED 200 0.32 0.76 5.95
1000 0.20 0.86 7.15
2000 0.29 0.78 6.19
SEED-IV 200 0.36 0.75 5.93
1000 0.19 0.88 7.24
2000 0.31 0.76 6.11
Localize-MI 200 0.15 0.92 8.37
1000 0.11 0.95 8.88
2000 0.11 0.95 8.94
Table 8: Impact of training timesteps on reconstruction quality
D.2.3 Cosine Schedule Offset Factor
In the cosine noise schedule for DDIM, the offset factor $s$ adjusts the smoothness and starting point of the noise variance curve to prevent instability from overly small initial noise levels. Concretely, $s$ introduces a phase shift in the cosine function, producing a more gradual noise increase at early timesteps—thereby avoiding abrupt noise jumps—while still covering the full variance range at later steps. Smaller values of $s$ yield gentler initial noise ramp-up, whereas larger $s$ accelerate early noise growth. Table 9 depicts the effect of cosine schedule offset factors on reconstruction quality.
Dataset $\boldsymbol{s}$ NMSE PCC SNR (dB)
SEED 0.005 0.20 0.86 7.15
0.010 0.24 0.84 7.02
0.025 0.26 0.83 6.93
SEED-IV 0.005 0.20 0.86 7.17
0.010 0.19 0.88 7.24
0.025 0.20 0.85 7.11
Localize-MI 0.005 0.11 0.95 8.94
0.010 0.15 0.93 8.41
0.025 0.18 0.91 8.23
Table 9: Effect of cosine schedule offset factor $s$ on reconstruction quality
Appendix E Downstream Task
E.1 Classification Feature Extraction
E.1.1 Differential Entropy Feature
Raw EEG at $1000$ Hz is downsampled to $200$ Hz, band-pass filtered (1501–50 Hz) with a $6$th-order Butterworth filter, and segmented into non-overlapping $1$ s windows ($200$ samples). Each window is transformed by STFT (Hanning window, $200$-point length, 256-point FFT). We compute band-specific power $E$ for $\delta\,(1\text{--}3\,\mathrm{Hz})$, $\theta\,(4\text{--}7\,\mathrm{Hz})$, $\alpha\,(8\text{--}13\,\mathrm{Hz})$, $\beta\,(14\text{--}30\,\mathrm{Hz})$, and $\gamma\,(31\text{--}50\,\mathrm{Hz})$ , normalize by the number of bins $N$, and define differential entropy (DE) feature as $\log(E/N)$ with a small constant added for numerical stability.
E.1.2 Power Spectral Density Feature
Power spectral density (PSD) feature features use the same STFT pipeline but report the mean squared magnitude (average power) in each band.
E.2 Random Forest Classifier
For emotion classification on SEED and SEED-IV and epileptic detection on Localize-MI, we employ a random forest with 100 trees. This set of hyperparameters balances nonlinearity modeling with computational efficiency, yielding robust performance on high-dimensional EEG features.
Appendix F Supplemental Ablation Results
Figure 10 presents additional ablation studies on SEED-IV and Localize-MI datasets, reporting NMSE, PCC, and SNR for the baseline LDM, LDM+SMM, LDM+RDM, and the full SRGDiff across various upsampling scales. These plots further illustrate the individual and combined contributions of our two modules to reconstruction quality.
Figure 10: Ablation results on SEED-IV and Localize-MI: comparison of NMSE, PCC, and SNR for LDM, LDM+SMM, LDM+RDM, and SRGDiff at 2×, 4×, and 8× upsampling and 2×, 4×, 8×, and 16× upsampling, respectively.
Table 10: Runtime (ms) of different methods on SEED, SEED-IV, and Localize-MI datasets.
Method SEED SEED-IV MI
SasDim 265.4 269.5 374.4
SADI 329.0 324.1 428.7
RDPI 318.1 315.9 421.2
DDPMEEG 549.3 558.2 841.7
ESTformer 3.51 3.55 5.03
STAD 232.6 227.8 385.9
Ours 63.0 62.1 92.5
Appendix G Reconstruction Visualization
Figure 11 through Figure 13 illustrate qualitative reconstructions on the three datasets. For each, we plot a single representative channel over time, comparing the ground‐truth high-density EEG (black) against STAD (blue), ESTformer (red), and SRGDiff (green). These overlays demonstrate SRGDiff’s closer alignment with the true waveform across diverse temporal patterns.
Figure 11: Reconstruction of a representative channel on Localize-MI (motor imagery): ground truth (black), STAD (blue), ESTformer (red), and SRGDiff (green).
Figure 12: Reconstruction of a representative channel on SEED-IV (emotion recognition): ground truth (black), STAD (blue), ESTformer (red), and SRGDiff (green).
Figure 13: Reconstruction of a representative channel on SEED (emotion recognition): ground truth (black), STAD (blue), ESTformer (red), and SRGDiff (green).